# Logistic Multidimensional Data Analysis for Ordinal Response Variables using a Cumulative Link function

### v1

@@ 2.3.2 CLMDU biplots
It follows that we can add circles to the biplot with center $\bm{v}_{r}$ and radius $a_{r_{c}}$, such that for points inside this circle the probability for responding higher than $c$ is larger than 0.5 while outside the circle this probability is smaller or equal to 0.5. Every variable point is therefore accompanied with $C_{r}-1$ circles representing the different probabilities. We illustrate the threshold circles in Figure 2b, where again $\bm{v}_{r}=[1,0.5]^{\prime}$ and $m_{r_{c}}=(-2.0,-1.5,-0.5)$. From the $m_{r_{c}}$ we obtain $a_{r_{c}}=(2.0,1.5,0.5)$. In some cases the radius can be negative, indicating that nowhere in the two-dimensional space a cumulative probability is larger than a half, and the circle is not drawn. In an usual biplot, several response variables are represented, so we have several points with surrounding circles jointly partitioning the two-dimensional space. Again each region in this space represents a particular response profile.
@@ 2.4 Restricted Models
In case we have predictor variables available for the observations, the principal scores or ideal points ($\bm{u}_{i}$) are defined to be linear combinations of the predictor variables, that is $\bm{u}_{i}=\bm{B}^{\prime}\bm{x}_{i}$. Numerical predictor variables can be included in the two-dimensional biplot as straight lines through the origin and with direction $b_{p2}/b_{p1}$. The positions of the observations can be obtained from the predictor variable axes by the process of interpolation, as outlined by Gower and Hand, (1996).
Categorical predictor variables, are recoded into dummy variables, where one of the categories is chosen as a reference category. In the biplot representation, we use points instead of variable axes for such predictor variables. The position of the reference category is in the origin of the low dimensional space, whereas the other categories are positioned at their corresponding estimates in $\bm{B}$.
For interpretation, the relationship between a predictor variable and a response variable is of interest. In the CLPCA biplots, for numerical predictor variales such a relationship is given by the angle between a predictor variable axis and that of a response variable. A sharp angle indicates a strong relationship, while an obtuse angle indicates the absence of a relationship. Furthermore, for every point along the predictor variable axis, a projection onto the response variable axis can be made to obtain a predicted value. For categorical predictor, the point representing a category can be projected onto the response variable axis can be made to obtain a predicted value.
For CLMDU biplots, the interpretation of predictor-response relationships is more involved, because these are single-peaked where with increasing values of a predictor first the response goes up and afterwards down again. Furthermore, although the effect of predictor variables is additive for obtaining the position of an observation (i.e., the ideal point), this additivity does not translate to the relationship towards the response variable.
@@ 3 Maximum Likelihood Estimation
In this section, we will develop an Expectation Majorization Minimization (EMM) algorithm to minimize the negative loglikelihood. We start by formulating the complete data negative loglikelihood, take the conditional expectation of this function in the E-step, and find a majorization function that can easily be minimized. The majorization function turns out to be a least squares function, so that in the inner loop of the algorithm, well known updating steps from least squares theory can be used.
@@ 3.1 Estimation of Structural part
$\mathbb{E}(\mathcal{L}_{c}(\theta_{ir}))\leq\mathbb{E}(\mathcal{L}_{c}(\tilde{\theta}_{ir}))+(\theta_{ir}-\tilde{\theta}_{ir})\mathbb{E}\left(\frac{\partial\mathcal{L}_{c}(\tilde{\theta}_{ir})}{\partial\theta_{ir}}\right)+\frac{1}{8}(\theta_{ir}-\tilde{\theta}_{ir})(\theta_{ir}-\tilde{\theta}_{ir})=\mathcal{M}(\theta_{ir}|\tilde{\theta}_{ir}).$
Let us define $p_{ir}=\frac{1}{1+\exp(-z_{ir}+\theta_{ir})}$ such that $\log f(z_{ir}-\theta_{ir})=\log p_{ir}(1-p_{ir})$. The partial derivative is
@@
Let $\lambda_{ir}=\tilde{\theta}_{ir}-4\xi_{ir}$ to obtain
@@
Now collecting all terms into a single function again, we obtain
$\mathbb{E}(\mathcal{L}_{c}(\bm{\theta}))\leq\mathcal{M}(\bm{\theta}|\tilde{\bm{\theta}})=\frac{1}{8}\sum_{i}\sum_{r}(\theta_{ir}-\lambda_{ir})^{2}+c_{2}.$
In the following four subsections, we work out this least squares loss function for the four different definitions of the structural part ($\theta_{ir}$).
@@ 3.1.1 PCA parametrisation of the Structural part
Remember that $\theta_{ir}=\langle\bm{u}_{i},\bm{v}_{r}\rangle$, so that the loss function in matrix algebra terms is
$\bm{\Lambda}=\bm{U}\bm{V}^{\prime},$
that is, we have to find a reduced rank approximation of the matrix $\bm{\Lambda}$ with elements $\lambda_{ir}$. Eckart and Young, (1936) showed that this can be done using a singular value decomposition
@@ 3.1.4 RMDU parametrisation of the Structural part
In the supervised analysis, we have that $\bm{U}=\bm{XB}$ and we need to estimate $\bm{B}$ instead of $\bm{U}$. Given the other parameters, the update of $\bm{B}$ is given as
@@ 3.2 Estimation of thresholds
To update the threshold parameters for each response variable we cannot use the complete data negative log-likelihood. However, we can use the default maximum likelihood estimator as in the proportional odds regression model. For this estimator $\bm{y}_{r}$ is the response variable and we use $\bm{\theta}_{r}$ as an offset without further predictor variables, so this gives estimates of the intercepts or thresholds. This provides estimates for response variable $r$. We repeat the procedure for each response variable.
@@ 3.3 Remarks on Algorithms
4 Application
We will use data from the International Social Survey Programme 2020, the Module on Environment, to illustrate our methods (ISSP Research Group,, 2022). The data we will use has $N=16465$ observations (7917 Male and 8548 Female) from 13 different countries and concerns attitudes and behavior related to environmental issues.
The thirteen countries are Austria (AT, 1112), Taiwan (TW, 1574), Finland (FI, 942), Germany (DE, 1365), Hungary (HU. 856), Iceland (IS, 867), Japan (JP, 964), New Zealand (NZ, 807), Philippines (PH, 1346), Russia (RU, 959), Slovenia (SL, 960), Switzerland (CH, 3650), and Thailand (TH, 1063). Also available are two background variables that will be used as predictors, Education was measured by the number of years education received, and working status with three classes.
The survey includes four pro-environmental behavior variables, measured on ordinal scales:
@@ 4 Application
Environmental concern was measured by the following statement: Generally speaking, how concerned are you about environmental issues, where participants could respond on a 5-point scale ranging from "Not at all concerned" (1) till "Very concerned" (5).
Environmental efficacy was measured by the following seven statements that each had a 5-point answer scale ranging from Agree Strongly (coded as 5) to Disagree Strongly (coded as 1):

### v2

@@ 2.3.2 CLMDU biplots
It follows that we can add circles to the biplot with center $\bm{v}_{r}$ and radius $a_{r_{c}}$, such that for points inside this circle the probability for responding higher than $c$ is larger than 0.5 while outside the circle this probability is smaller or equal to 0.5. Every variable point is therefore accompanied with $C_{r}-1$ circles representing the different probabilities. We illustrate the threshold circles in Figure 2b, where again $\bm{v}_{r}=[1,0.5]^{\prime}$ and $m_{r_{c}}=(-2.0,-1.5,-0.5)$. From the $m_{r_{c}}$ we obtain $a_{r_{c}}=(2.0,1.5,0.5)$. We see again that the two dimensional space is partitioned in several areas. These areas are now defined by the circles. Withing the smallest circle around the response variable, the participants are predicted to scores highest. Around the inner circle, we have regions in the form of tyres. Within such a band, participants are predicted to have the same score on the response variable. In some cases the radius can be negative (when $m_{r_{c}}$ is positive), indicating that nowhere in the two-dimensional space a cumulative probability is larger than a half, and the circle is not drawn.
In an usual biplot, several response variables are represented. Each of the points representing a response variable is surrounded by a set of circles. Circles of different response variables cross, creating regions that each represent a particular response profile.
@@ 2.4 Restricted Models
In case we have predictor variables available for the observations, the principal scores or ideal points ($\bm{u}_{i}$) are defined to be linear combinations of the predictor variables, that is $\bm{u}_{i}=\bm{B}^{\prime}\bm{x}_{i}$. To include the predictor variables in the biplot, we distinguish between numerical and categorical predictor variables. Numerical predictor variables can be included in the two-dimensional biplot as straight lines through the origin and with direction $b_{p2}/b_{p1}$. The positions of the observations can be obtained from the predictor variable axes by the process of interpolation, as outlined by Gower and Hand, (1996). The interpolation process is similar to vector addition, that is, we create vectors starting in the origin and along the variable axes to the observed predictor value for each predictor variables. To obtain the position of the participant, we have to add these vectors.
Categorical predictor variables, are recoded into dummy variables, where one of the categories is chosen as a reference category. In the biplot representation, we use points instead of variable axes for such predictor variables. The position of the reference category is in the origin of the low dimensional space, whereas the other categories are positioned at their corresponding estimates in $\bm{B}$. We can consider categorical predictor variables as “jumps”. When the categorical predictor is equal to the reference category, no jump is made. When the categorical predictor variable is equal to another category, a jump from one position to another is made. This jump changes the location of variable axes for the numerical predictor variables.
Variable axes for the numerical predictor variables and points for the categorical predictor variables are added to the biplots for CLPCA or CLMDU. For interpretation, the relationship between a predictor variable and a response variable is of interest. In the CLRRR biplots, for numerical predictor variables such a relationship is given by the angle between a predictor variable axis and that of a response variable (see Section 2.1). A sharp angle indicates a strong relationship, while an obtuse angle indicates the absence of a relationship. Furthermore, for every point along the predictor variable axis, a projection onto the response variable axis can be made to obtain a predicted value. For categorical predictor, the point representing a category can be projected onto the response variable axis can be made to obtain a predicted value.
For CLRMDU biplots, the interpretation of predictor-response relationships is more involved, because these are single-peaked where with increasing values of a predictor first the response goes up and afterwards down again. Furthermore, although the effect of predictor variables is additive for obtaining the position of an observation (i.e., the ideal point $\bm{u}_{i}$), this additivity does not translate to the relationship towards the response variable.
@@ 3 Maximum Likelihood Estimation
In this section, we will develop an Expectation Majorization Minimization (EMM) algorithm to minimize the negative loglikelihood. This algorithm is a combination of the EM-algorithm often used for latent variable models (McLachlan and Krishnan,, 2007) and the MM-algorithm (Hunter and Lange,, 2004; Heiser,, 1995).
We start by formulating the complete data negative loglikelihood, take the conditional expectation of this function in the E-step, and find a majorization function that can easily be minimized. The majorization function turns out to be a least squares function, so that in the inner loop of the algorithm, well known updating steps from least squares theory can be used.
@@ 3.1 Estimation of Structural part
$\mathbb{E}(\mathcal{L}_{c}(\theta_{ir}))\leq\mathbb{E}(\mathcal{L}_{c}(\tilde{\theta}_{ir}))+(\theta_{ir}-\tilde{\theta}_{ir})\mathbb{E}\left(\frac{\partial\mathcal{L}_{c}(\tilde{\theta}_{ir})}{\partial\theta_{ir}}\right)+\frac{1}{8}(\theta_{ir}-\tilde{\theta}_{ir})(\theta_{ir}-\tilde{\theta}_{ir})=\mathcal{M}(\theta_{ir}|\tilde{\theta}_{ir}),$
where $\mathcal{M}(\theta_{ir}|\tilde{\theta}_{ir})$ is the majorization function. Let us define $p_{ir}=\frac{1}{1+\exp(-z_{ir}+\theta_{ir})}$ such that $\log f(z_{ir}-\theta_{ir})=\log p_{ir}(1-p_{ir})$. The partial derivative is
@@
Now, let us define, what we will call working responses
$\lambda_{ir}=\tilde{\theta}_{ir}-4\xi_{ir}$ (2)
to obtain
@@
Now collecting all terms into a single function, we obtain
$\mathbb{E}(\mathcal{L}_{c}(\bm{\theta}))\leq\mathcal{M}(\bm{\theta}|\tilde{\bm{\theta}})=\frac{1}{8}\sum_{i}\sum_{r}(\theta_{ir}-\lambda_{ir})^{2}+c_{2},$
a least squares function with the working responses $\lambda_{ir}$. In the following four subsections, we work out this least squares loss function for the four different definitions of the structural part ($\theta_{ir}$).
@@ 3.1.1 PCA parametrisation of the Structural part
Remember that $\theta_{ir}=\langle\bm{u}_{i},\bm{v}_{r}\rangle$, so that the loss function equals
$\sum_{i}\sum_{r}(\lambda_{ir}-\bm{u}_{i}^{\prime}\bm{v}_{r})^{2},$
which can be written in matrix algebra terms as
$\|\bm{\Lambda}-\bm{U}\bm{V}^{\prime}\|^{2}.$
We have to find a reduced rank approximation of the matrix $\bm{\Lambda}$ with elements $\lambda_{ir}$. Eckart and Young, (1936) showed that this can be done using a singular value decomposition
@@ 3.1.4 RMDU parametrisation of the Structural part
When predictor variables are available, we constrain $\bm{U}=\bm{XB}$ and we need to estimate $\bm{B}$ instead of $\bm{U}$. We can use the algorithm of MDU described in the previous section and replace the updating equation for $\bm{U}$ (i.e., Equation 9) with an updating equation for $\bm{B}$, that is
@@ 3.2 Estimation of thresholds
To update the threshold parameters for each response variable we cannot use the complete data negative log-likelihood. However, we can use the default maximum likelihood estimator as in the proportional odds regression model. For this estimator $\bm{y}_{r}$ is the response variable and we use $\bm{\theta}_{r}$ as an offset (i.e., a predictor variable with regression weight fixed to 1) without any further predictor variables. This gives maximum likelihood estimates of the intercepts or thresholds for response variable $r$. We repeat the procedure for each response variable.
@@ 3.3 Remarks on Algorithms
3.4 Algorithm schemes
To summarize, we give algorithm schemes in four algorithm boxes. Algorithm 1 shows the procedure for CLPCA, algorithm 2 for CLRRR, algorithm 3 for CLMDU, and algorithm 4 for CLRMDU.
Algorithm 1 CLPCA algorithm.
1:procedure clpca($\bm{Y},\bm{m},\bm{U},\bm{V}$)
2:     Predefine: maxouter, $\epsilon$
3:     Compute structural part $\bm{\theta}$
4:     Assess $\mathcal{L}_{o}^{0}(\bm{\theta},\bf{m})$
5:     for $t_{1}\leftarrow 1,\text{maxouter}$ do
6:         Compute working responses $\bm{\Lambda}=\{\lambda_{ir}\}$ using equation (2)
7:         SVD of $\bm{\Lambda}$
8:         Update $\bm{U}$ using equation (3)
9:         Update $\bm{V}$ using equation (4)
10:         Update $\bm{m}$ by directly maximizing the likelihood
11:         if $\mathcal{L}_{o}^{t_{1}}(\bm{\theta}\bm{m})-\mathcal{L}_{o}^{t_{1}-1}(\bm{\theta},\bm{m})<\epsilon$: break
12:     end for
13:     return($\bm{m},\bm{U},\bm{V}$)
14:end procedure
Algorithm 2 CLRRR algorithm.
1:procedure clrrr($\bm{Y},\bm{X},\bm{m},\bm{B},\bm{V}$)
2:     Predefine: maxouter, $\epsilon$
3:     Compute structural part $\bm{\theta}$
4:     Assess $\mathcal{L}_{o}^{0}(\bm{\theta},\bf{m})$
5:     for $t_{1}\leftarrow 1,\text{maxouter}$ do
6:         Compute working responses $\bm{\Lambda}=\{\lambda_{ir}\}$ using equation (2)
7:         GSVD of $\bm{\Lambda}$
8:         Update $\bm{B}$ using equation (5)
9:         Update $\bm{V}$ using equation (6)
10:         Update $\bm{m}$ by directly maximizing the likelihood
11:         if $\mathcal{L}_{o}^{t_{1}}(\bm{\theta}\bm{m})-\mathcal{L}_{o}^{t_{1}-1}(\bm{\theta},\bm{m})<\epsilon$: break
12:     end for
13:     return($\bm{m},\bm{B},\bm{V}$)
14:end procedure
Algorithm 3 CLMDU algorithm.
1:procedure clmru($\bm{Y},\bm{m},\bm{U},\bm{V}$)
2:     Predefine: maxouter, maxinner, $\epsilon_{1}$, $\epsilon_{2}$
3:     Compute structural part $\bm{\theta}$
4:     Assess $\mathcal{L}_{o}^{0}(\bm{\theta},\bf{m})$
5:     for $t_{1}\leftarrow 1,\text{maxouter}$ do
6:         Compute working responses $\bm{\Lambda}=\{\lambda_{ir}\}$ using equation (2)
7:         Assess $\mathcal{M}^{0}(\bm{U},\bf{V})$
8:         for $t_{2}\leftarrow 1,\text{maxinner}$ do
9:              Update $\bm{U}$ using equation (9)
10:              Update $\bm{V}$ using equation (10)
11:              Assess $\mathcal{M}^{t2}(\bm{U},\bf{V})$
12:              if $\mathcal{M}^{t_{2}}(\bm{U},\bm{V})-\mathcal{M}^{t_{2}-1}(\bm{U},\bm{V})<\epsilon_{2}$: break
13:         end for
14:         Update $\bm{m}$ by directly maximizing the likelihood
15:         if $\mathcal{L}_{o}^{t_{1}}(\bm{\theta}\bm{m})-\mathcal{L}_{o}^{t_{1}-1}(\bm{\theta},\bm{m})<\epsilon_{1}$: break
16:     end for
17:     return($\bm{m},\bm{U},\bm{V}$)
18:end procedure
Algorithm 4 CLRMDU algorithm.
1:procedure clrmdu($\bm{Y},\bm{X},\bm{m},\bm{B},\bm{V}$)
2:     predefine: maxouter, maxinner, $\epsilon_{1}$, $\epsilon_{2}$
3:     Compute structural part $\bm{\theta}$
4:     Assess $\mathcal{L}_{o}^{0}(\bm{\theta},\bf{m})$
5:     for $t_{1}\leftarrow 1,\text{maxouter}$ do
6:         Compute working responses $\bm{\Lambda}=\{\lambda_{ir}\}$ using equation (2)
7:         Assess $\mathcal{M}^{0}(\bm{U},\bf{V})$
8:         for $t_{2}\leftarrow 1,\text{maxinner}$ do
9:              Update $\bm{B}$ using equation (11)
10:              Update $\bm{V}$ using equation (10)
11:              Assess $\mathcal{M}^{t2}(\bm{U},\bf{V})$
12:              if $\mathcal{M}^{t_{2}}(\bm{U},\bm{V})-\mathcal{M}^{t_{2}-1}(\bm{U},\bm{V})<\epsilon_{2}$: break
13:         end for
14:         Update $\bm{m}$ by directly maximizing the likelihood
15:         if $\mathcal{L}_{o}^{t_{1}}(\bm{\theta}\bm{m})-\mathcal{L}_{o}^{t_{1}-1}(\bm{\theta},\bm{m})<\epsilon_{1}$: break
16:     end for
17:     return($\bm{m},\bm{B},\bm{V}$)
18:end procedure
4 Applications
In this section, we will show three applications of our modeling framework. In the first application, we analyze part of the data described by Fabbricatore et al., (2024) about students responses to exam questions about statistics and relationships with several psychological variables like attitude and anxiety. We will analyze these data using the cumulative logistic reduced rank model. In the second application, we use data from the International Social Survey program to investigate the relationship between environmental attitudes and pro-environmental behaviour. In this analysis, we use both the dominance and proximity models to highlight the differences. Finally, we show a third application again using data from the International Social Survey program. We use data from seven Likert scales to investigate differences between countries in environmental efficacy. We use the cumulative logistic restricted multidimensional unfolding model for this analysis.
4.1 Students’ Performance for Statistical Tests
The subset of data we use for this analysis involves 138 university students and their responses to ten questions in an exam about the application of statistics to certain topics. The ten items are described in detail in Appendix A. Each response is coded as wrong (0), partially correct (1), or correct (2), that is, as a three categories ordinal response.
Prior to the start of the courses on statistics, several psychological tests were conducted. Besides their answers on the ten items, for each student information is available on their gender (1 = female), age in years, and several measurements of mathematical knowledge and psychological factors, each assessed through validated psychometric instruments. The variables that we use are:
  • Mathematical knowledge measured using the Mathematical Prerequisites for Psychometrics scale (PMP);
  • Statistical anxiety with three scale scores referring to examination anxiety (SASa), interpretation anxiety (SASi), and fear of asking for help (SASf);
  • Attitudes toward Statistics with four scales: affect (SATSa), cognitive competence (SATSc), value (SATSv), and difficulty (SATSd);
  • Motivated strategy for learning with four scales referring to self-efficacy (MSLe), test anxiety (MSLt), cognitive strategies (MSLc), and self-regulation (MSLsr);
  • Academic procrastination measured using a single scale (APS);
  • Academic motivation also measured using a single scale (AMS);
  • Student engagement in statistics measured through three scales: affective engagement (ENGa), behavioural engagement (ENGb), and cognitive engagement (ENGc).
Psychometric properties of these scales are all satisfactory to good, see Fabbricatore et al., (2024) for details, where also more detailed descriptions of the scales can be found. The scores on these scales serve, together with gender and age, as predictor variables. The responses to the ten statistical application items serve as response variables.
As there is no prior information whether the 10 items comprise a unidimensional or multidimensional construct, we will start fitting models in one till three dimensions and select an optimal dimensionality. Subsequently, we will verify which of the predictor variables influence the responses. Fit statistics for the one-, two-, and three dimensional models are shown in Appendix A, where it can be seen that the two-dimensional model has the lowest AIC, while the lowest BIC is obtained for the unidimensional model. We proceed with the two-dimensional model. Leaving out each (set of) predictor variables, we obtain the fit statistics in the lower part of Table A1, showing that Age, Statistical Anxiety, the Motivated Strategies for Learning Questionnaire, and the Academic Motivation Scale can be left out without significant loss of fit (AIC based conclusion).
Figure 3: Estimated configuration for students data. The dark green lines represent the response variables, the blue lines represent the predictor variables. Variable labels are placed on the positive side of the variables, that are the sides with the largest values. In the upper right corner the labels of ENGa and A1d overlap.
The final biplot is shown in Figure 3. The implied coefficients are shown in Table A2 (Appendix A). The figure and table can be used together to come to a interpretation of the final model. Let us first look at the configuration of the response variables (items A1a, …, A2g) represented by the green solid lines with markers 0|1 and 1|2. By inspecting the direction of the variable axes, we see most responses are strongly correlated, that is, the variable axes have small angles and all point to the right hand side of the figure. A notable exception is item A2g, and to a smaller extend A1d. The general pattern is that students on the right hand side of the figure correctly answer the items, while students on the left hand side often make mistakes. For items A2g and A1d, students in the top of the configuration make the items correct, while at the bottom of the figure they tend to fail. Each variable axis also has two markers (0|1 and 1|2) indicating the difficulty of the item. For response variable A2a, for example, we see that these two markers are strongly to the left. Projecting the students on the axes we see that most fall in the region of a correct answer. This item is relatively easy. For item A2i, however, the marker 1|2 is far to the right and the projection of student points on this axis is rarely in the region of a correct answer. Note that item A2i is also the item with the least correct answers in the data, whereas A2a has many correct responses.
Now let us further inspect the predictor variables (indicated by the light blue variable axes) and the positions of the students in the configuration. The predictor variables are represented with a variable axis that has a solid part and a dotted part: the solid part indicates the observed range in the data, from the minimum to the maximum observed value, and as such functions as a measure of effect size. The dotted part only extends the variable axis to the edge of the biplot where the variable label is placed. It can be seen that PMP has a long solid variable axis, that is, mathematical knowledge prior to the courses makes large differences. Students who score high on mathematics (PMP) are positioned on the right hand side. It follows that a high PMP score is predictive for answering most items correctly. Similarly, high behavioural engagement (ENGb) and a strong positive feeling about statistics (i.e., a high score on the affect scale SATSa) result in positions more to the right hand side of the biplot, where the model indicates correct responses to the items.
The predictor variable affective engagement (ENGa) points to the upper right corner, so is a good predictor for a correct response on A1d and A2g. Students who consider statistics to be a difficult topic (SATSd), in contrast, will be on the lower left corner, that is, the variable axis points in the opposite direction. Therefore, lower scores on this variable predict a correct response on A1d and A2g.
Higher scores on academic procrastination (APS) and the value and cognitive competence scales of attitudes towards statistics (SATSv and SATSc, respectively) result in positions more to the top left. These points project low on most of the response variables, indicating wrong answers, but high (i.e., correct answers) on A2g and A1d. In contrast, higher scores on cognitive engagement (ENGc) result in positioning more towards the bottom right, an opposite response pattern compared to academic procrastination (APS) and the two attitudes scales (SATSv and SATSc). Finally, the gender variable (G) points towards the left: When a boy and a girl have equal values for the other predictor variables, girls are positioned more to the left hand side, indicating less favorable answers to the ten statistics items.
4.2 ISSP data: Pro-environmental behaviour
In this application and the next, we use from the International Social Survey Programme 2020, the Module on Environment (ISSP Research Group,, 2022). In this first analysis, we focus on the data from Thailand ($N=1063$) and the responses to four pro-environmental behaviour variables, measured on ordinal scales:
@@ 4.2 ISSP data: Pro-environmental behaviour
We are interested in the relationship between environmental concern and environmental efficacy and these four behavioural items. Environmental concern (EC) was measured by the following statement: Generally speaking, how concerned are you about environmental issues, where participants could respond on a 5-point scale ranging from "Not at all concerned" (1) till "Very concerned" (5). Environmental efficacy (EE) was measured by averaging responses to seven statements that each had a 5-point answer scale ranging from Agree Strongly (5) to Disagree Strongly (1). We also use gender, age, and number of years of education as predictor variables in the analysis. In psychological research, the relationship between attitudes and behaviour is often of interest. Therefore, our primary focus is on the relationship between environmental concern and efficacy and the four behavioural items, controlling for the other variables. We will analyse the data using both the dominance and proximity perspective and contrast the two analyses.
In Table 1, we show the fit statistics of both analyses in dimensionalities 1, 2, and 3. The AIC and BIC disagree on the optimal dimensionality, that is, AIC points to the three dimensional solutions, while BIC points towards the two-dimensional models. We will focus on the more parsimonious two-dimensional solutions. Comparing the dominance (CLRRR) and proximity (CLRMDU) perspective, it can be seen that the fit statistics for the proximity model are a little better with lower deviance and AIC.
We will inspect and interpret the visualization for the proximity model, the solution of the dominance model can be found in Appendix B (we will comment on it at the end of this section). The biplot is shown in Figure 4, where it can be seen that the position of three of the four response variables is on the outside of the configuration, while one is more centrally located (i.e., OUT). This means that, while the predictors have a dominance relationship with the response variables MEAT, RECYCLE, and AVOID, they have a proximity relationship with OUT.
CLRRR CLRMDU
dimensionality deviance AIC BIC deviance AIC BIC
1 11826.94 11876.94 12001.16 11839.25 11891.25 12020.44
2 11777.22 11839.22 11993.25 11756.83 11824.83 11993.77
3 11754.77 11824.77 11998.68 11733.53 11815.53 12019.25
Table 1: Fit statistics for the dominance (CLRRR) and proximity (CLRMDU) analysis of pro-environmental behaviour in one, two, and three dimensions.
Let us look at the biplot in more detail. The two predictor variables of interest, environmental concern (EC) and efficacy (EE), are represented by variable axes that run from the right hand side to the left hand side, meaning that persons that score high on these variables are located at the left hand side of the biplot, while persons that score low on these variables are positioned at the right hand side in the biplot. The control variables, education and age also run from right (lower scores) to left (higher scores). Females are located more to the top left compared to males. The higher a participant scores on EE or EC the closer they get to the response variables RECYCLE and AVOID, meaning that the higher a participant scores on the two attitude variables the more pro-environmental behaviour is reported. For these two variables only the radii for never | sometimes (1|2) and sometimes | often (2|3) are positive, meaning that participants are never classified in the response category always. The biplots including the decision regions per response variable are shown in Figure B1 in Appendix B. These two response variables follow a dominance response process because all participant are located on one side of the point representing the response variable.
Figure 4: Biplot for the cumulative logistic restricted multidimensional unfolding solution relating environmental attitudes with pro-environmental behaviour.
The response variable OUT lies close to the origin of the two-dimensional solution. The radii for boundaries never | several times per year (1|2) and several times per year | several times per month (2|3) are positive. No participant will be classified in the class of daily (5) or several times per week (4) outside activities, as the estimated radius for these boundaries is negative. Note that younger participants are closer to the point for this response variable, indicating younger people more often engage in outside activities. With increasing scores on environmental efficacy (and average values for the other variables) the behaviour changes from several times per year, to several times per month, to several times per year, and extrapolating to never, a single peaked response pattern. A similar response function can be obtained for the predictor environmental concern (EC).
Finally, we inspect the response variable MEAT. All radii are positive and therefore participants can be classified in each of the categories. The classification regions are shown in Figure B1, where for this data set the participants fall in 0, 1, 2, and 3 days per week eating meet (notice the reverse where not eating meat (0 days) is coded as 8 indicating the category with most pro-environmental behaviour). No participant is classified as eating meat more days per week, these regions fall on the lower bottom of the figure, where no participants are located. We see that some participants will be classified as typically eating no meat (within the smallest circle around MEAT). There is not a single predictor pointing in this direction, but a combination of female, higher age, relatively low EC, and a low number of years of education will result in this classification.
As pointed out before, the answers to three out of the four response variables follow a dominance pattern, whereas only the responses to OUT follow a proximity pattern. This response variable (OUT) is probably the reason that the proximity model fits these data better. The biplot for the two-dimensional dominance model is shown as Figure B2, where the interpretation for MEAT, AVOID and RECYCLE closely follows the interpretation given here. The direction of the response variables is very similar to the positions in Figure 4. The pattern for response variable OUT is different, now being monotonic and negatively related with environmental efficacy, age, and being female.
4.3 ISSP data: Environmental Efficacy
In this second analysis using the ISSP data, we focus on differences between countries concerning environmental efficacy, controlling for gender, education and age. The seven items related to environmental efficacy are: